OpenStudy (.whitedragon.):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3). (Points : 1) y = 7x - 10 y = -1/7 x + 2 y = 7x - 3 y = -7x + 10

OpenStudy (igreen):

Parallel lines have the same slope..so what's the slope of the second line going to be?

OpenStudy (.whitedragon.):

htfsik? idk

OpenStudy (kamibug):

Slope-intercept form is y = mx + b where m is the slope. When looking at the equation you are given, y = 7x +2, what is the m? :)

OpenStudy (igreen):

^^ It's in the form of y = mx + b, where m = slope.

OpenStudy (.whitedragon.):

7

OpenStudy (kamibug):

Yup, and like @iGreen said, parallel lines have the same slope. :) So which answers can you cross out?

OpenStudy (.whitedragon.):

so c ?

OpenStudy (igreen):

Well, there are only two answers we have, which are: A. y = 7x - 10 and C. y = 7x - 3 Now we have to find out which one crosses through (1, -3). To do this, we can plug in (1, -3) into both equations and see which one fits. y = 7x - 10 -3 = 7(1) - 10 -3 = 7 - 10 What's 7 - 10? @.WhiteDragon.

OpenStudy (igreen):

@.WhiteDragon. Are you there?

OpenStudy (.whitedragon.):

sorry yes

OpenStudy (.whitedragon.):

3 so c is the answer ?

OpenStudy (igreen):

No. Check again, what's 7 - 10?

OpenStudy (kamibug):

In slope-intercept form the "b" is the y-intercept. In the coordinate of a y-intercept, the x value always equals 0 because it's where the line crosses the y-axis. So another way for us to find what that value of "b" is, is to plug in the into we know into the slope formula and solve for that other y. So m is 7, and one of the y's is -3, and x2 is 1 and x1 is going to be 0 because we wanna find y when it's the intercept. So we can use this equation and solve for y1...